OpenStudy (anonymous):

Suppose that dy/dx=(x-y)^2+1. Anyone know how to get the general solution to the differential equation?

OpenStudy (anonymous):

x-y = t 1 - dy/dx = dt/dx so it becomes 1-dt/dx = t^2 +1

OpenStudy (anonymous):

get it?????????

OpenStudy (anonymous):

Yeah, thanks